Put into standard form:
x-2y+7=0 (-4,0)< x,y
I've tried to get y by itself:

y= one-half x-3.5 > 0=one-half(-4)-3.5 > 0=-2-3.5 > 0=-5.5

I know that that isn't right but I can't figure out what I did wrong. Any help would be greatly appreciated.

Answer by longjonsilver(2297)

x-2y+7 = 0
x+7 = 2y
2y = x+7
y = (1/2)x+(7/2)

so you are correct upto here.

To be honest, your question says get it into standard form... you just did. So end of question.

I have no idea why you quoted the (-4,0) point. Unless you want the line parallel to the quoted one that passes through point (-4,0). If this is so, then you need to say so.

Its equation will be of the form y=mx+c. If parallel, then the gradient will be the same, namely (1/2), so that

y=(1/2)x + c. Now to find c by putting in the x and y values:
0=(1/2)(-4) + c
0= -2 + c
--> c = 2

so equation is y = (1/2)x + 2

jon.